# Quantum kinetic approach to the Schwinger production of scalar particles in an expanding universe

### v2

@@ Abstract
We study the Schwinger pair creation of scalar charged particles by a homogeneous electric field in an expanding universe in the quantum kinetic approach. We introduce an adiabatic vacuum for the scalar field based on the Wentzel–Kramers–Brillouin solution to the mode equation in conformal time and apply the formalism of Bogolyubov coefficients to derive a system of quantum Vlasov equations for three real kinetic functions. Compared to the analogous system of equations previously reported in the literature, the new one has two advantages. First, its solutions exhibit a faster decrease at large momenta which makes it more suitable for numerical computations. Second, it predicts no particle creation in the case of conformally coupled massless scalar field in the vanishing electric field, i.e., it respects the conformal symmetry of the system. We identify the ultraviolet divergences in the electric current and energy-momentum tensor of produced particles and introduce the corresponding counterterms in order to cancel them.
@@ 1 Introduction
One of the most fascinating phenomena occurring in a strong electromagnetic field is the Schwinger effect or the creation of particle–antiparticle pairs from physical vacuum [1, 2, 3]. It was predicted almost a hundred years ago and until the present day has not been observed in the laboratory because of extremely high value of the required electric field $E\sim 10^{18}\,\text{V}/\text{m}$ [4]. Nevertheless, such enormous fields might exist in the universe, e.g., around the compact objects like neutron stars or black holes [5, 6] or during the early stages of evolution (see, e.g., Refs. [7, 8, 9]). The Schwinger effect in the early universe is often considered within the context of inflationary magnetogenesis models which include the coupling of the electromagnetic field to the inflaton or to the spacetime curvature [10, 11, 12]. The simplest scenario for the Schwinger pair production during inflation which can be studied analytically is the case of a constant and homogeneous electric field (with possible presence of a collinear magnetic field) in de Sitter spacetime which has been studied in much detail in the literature [7, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22]. The results of this simple approach has been used for a lot of phenomenological applications [23, 8, 24, 25, 26, 27, 28, 29, 9, 30, 31, 32, 33]; however, in realistic inflationary models both underlying assumptions, the constant in time gauge field and purely de Sitter expansion of the universe, typically do hot hold. In a time-dependent electromagnetic field, the Schwinger pair production has a more complicated dynamics. This was demonstrated by using the kinetic approach both in Minkowski space [34, 35, 36, 37, 38, 39, 40, 41] and in the expanding universe [42, 43]. In contrast to the case of constant field, the Schwinger induced current exhibits the non-Markovian character depending not only on the field at a given moment of time but also on the prehistory. Moreover, due to inertial properties of charge carriers the current is retarded with respect to the changes in the electric field which leads to the oscillatory behavior of the induced current and the electric field. In the kinetic approach, the particle creation is described by the Schwinger source term which can be constructed phenomenologically [34, 35, 42] or derived from the first principles [36, 37, 38, 39, 40, 41, 43]. The latter way, although being much more involved, allows to capture the effects of particle statistics and the nonlocal in time character of the pair-creation process. There is, however, a fundamental issue which makes the application of such a first-principles approach ambiguous in the case of the expanding universe. This is a well-known in the literature problem of defining the vacuum state and particles in nonstationary backgrounds which do not match the Minkowski spacetime in the asymptotic past and future [44, 45, 46, 47]. The result for the number of produced particles depends on the definition of physical vacuum state or, in other words, on the choice of the observer. In this work, we revisit the problem of pair creation of the scalar particles by a time-dependent electric field in the expanding Friedmann-Lemaître-Robertson-Walker (FLRW) universe previously considered in Ref. [43]. We mostly follow the same strategy in order to derive the system of quantum kinetic equations describing the Schwinger pair production in this system. However, in the present study, we perform all computations in terms of the conformal time rather then physical time considered in Ref. [43]. Despite the seemingly trivial change, we find that the definition of the adiabatic vacuum is different compared to the previous work and, consequently, the final system of equations appears to be different. We analyze the ultraviolet (UV) behavior of its solutions and show that kinetic functions exhibit faster decrease at large momenta as compared to the system in Ref. [43] which makes the new system of equations more attractive for numerical analysis. Also the new setup respects the conformal symmetry of the problem and does not lead to the particle production in the case of conformally coupled massless scalar field in the absence of the electric field. The aim of this work is not to show that the computations in Ref. [43] are incorrect, but to demonstrate that there is a different and more convenient way to derive the system of quantum kinetic equations. Although we expect that the numerical result in the case of a strong electric field $|e\boldsymbol{E}|\gg H^{2},\,m^{2}$ (where $e$ is the electric charge, $H$ is the Hubble parameter, and $m$ is the particle’s mass) would be the same for both approaches, there might be some numerical differences in the case of a weak field where the particle production by the electric field and the time-varying metric are comparable. The rest of the article has the following structure. In Sec. 2 we consider the evolution of a quantum scalar field on a classical background of the electric field and the expanding universe. In particular, we define the adiabatic vacuum and use the formalism of the Bogolyubov coefficients in order to derive the system of quantum kinetic equations. Then, in Sec. 3 we determine the basic observables – electric current and energy-momentum tensor of the produced particles – and express them in terms of the kinetic functions. In Sec. 4, we realize the renormalization program in order to separate the UV divergent contributions to the observables and cancel them by introducing the counterterms to the action. Section 5 is devoted to conclusions. In Appendix A, we determine the asymptotical behavior of the kinetic functions in the limit of large momenta. Throughout the work, we use natural units and set $\hbar=c=1$. We assume that the universe is described by a spatially flat FLRW metric. In terms of conformal time $\eta$ it has the form $g_{\mu\nu}=\operatorname{diag}(a^{2},\,-a^{2},\,-a^{2},\,-a^{2})$, where $a=a(\eta)$ is the scale factor.
@@ 2 Quantum Vlasov equations for the scalar field
where the term $\mathcal{L}_{\mathrm{int}}$ describes the coupling of the electromagnetic field to some generic field $\phi$. Although this interaction is necessary to produce sufficiently strong electromagnetic field needed for the Schwinger pair production, in this work we will not stick to any specific model and will assume that the electromagnetic field already exists in the Universe. Moreover, we consider the following configuration of the gauge field: (i) the magnetic field $\boldsymbol{B}$ is negligibly small compared to the electric one $\boldsymbol{E}$; (ii) the coherence length of the electric field $\lambda_{E}$ is much greater than any other physically relevant length scale in the problem, e.g., the Schwinger pair production scale $l_{\mathrm{S}}\sim|e\boldsymbol{E}|^{-1/2}$ or the Hubble scale $l_{H}\sim H^{-1}$. The latter condition allows us to treat the electric field spatially homogeneous and dependent only on time $\boldsymbol{E}=\boldsymbol{E}(\eta)$. Such a field configuration can be described by the vector potential (in the Coulomb gauge) $A_{\mu}=(0,\,\boldsymbol{A}(\eta))$ with $\boldsymbol{A}^{\prime}(\eta)=-a^{2}\boldsymbol{E}$. With the factor $a^{2}$, the electric field $\boldsymbol{E}$ is a physical field measured by the comoving observer. Finally, we go back to the last term in Eq. (1) which represents the Lagrangian density of the complex scalar charged field $\chi$ with mass $m$ and charge $e$:
@@
Here $\lvert\chi\rvert^{2}\equiv\chi^{\dagger}\chi$, $\mathcal{D}_{\mu}=\partial_{\mu}-ieA_{\mu}$ is the covariant derivative acting on the scalar field, and $\xi$ is the coupling constant responsible for the nonminimal coupling of the scalar field $\chi$ to gravity. Varying action in Eq. (1) with respect to $\chi^{\dagger}$, we obtain the equation of motion for the scalar field $\chi$:
@@
The factor $a(\eta)$ in denominator in Eq. (6) was introduced for further convenience. Inserting the expansion (6) into the equation of motion (5) we get the mode equation governing the evolution of the mode function $\chi_{\boldsymbol{k}}(\eta)$:
@@
It has an oscillator-like form with the time-dependent effective frequency
@@
It is easy to observe from Eq. (9) that in the absence of an electromagnetic field in a non-expanding Universe, the frequency $\Omega_{\boldsymbol{k}}$ does not depend on time $\eta$. Then, Eq. (8) has two independent solutions which describe the positive- and negative-frequency modes. If the initial condition imposed at a certain moment of time corresponds, e.g., to a positive-frequency mode, the solution remains to be positive-frequency all the time. Therefore, we conclude that the creation of particles in Minkowski spacetime does not occur. Moreover, even in the expanding universe, for the massless scalar particle, $m=0$, with the conformal nonminimal coupling to gravity, $\xi=1/6$, and in the absence of the electric field, equation of motion (8) has a very simple oscillator-like form $\chi^{\prime\prime}_{\boldsymbol{k}}+k^{2}\chi_{\boldsymbol{k}}=0$ whose positive-frequency solution is a well-known Bunch-Davies vacuum [48]:
@@
Being prepared in this state initially, the scalar field will remain in it forever meaning that there is no particle production even in the expanding universe. This is a consequence of the conformal invariance of the action of a massless conformally-coupled scalar field [47]. Here, we would like to note that this conformal symmetry becomes explicit only if we are working in the FLRW metric written in terms of the conformal time which is conformally flat. The situation changes in the presence of a nonzero electric field in the expanding universe. Indeed, the value of $\Omega_{\boldsymbol{k}}$ becomes dependent on the conformal time $\eta$, and therefore, in general, it is impossible to find exact general solution to the mode equation (8) and separate positive- and negative-frequency modes. Nevertheless, one can still construct an approximate solution by employing the Wentzel–Kramers–Brillouin approximation up to a certain adiabatic order. For example, in the zeroth adiabatic order, the positive-frequency solution has the form
@@
and at the moment of time $\eta_{0}$ they satisfy the initial conditions $\alpha_{\boldsymbol{k}}(\eta_{0})=1$ and $\beta_{\boldsymbol{k}}(\eta_{0})=0$ (by construction of the adiabatic vacuum at the moment $\eta_{0}$). It is straightforward to show that the mode equation (8) is identically satisfied if the evolution of the Bogolyubov coefficients is determined by the following system of equations:
@@
The main advantages of this system of equations compared to (2) is that it is real and does not contain fast oscillating coefficients. In order to rewrite it in the final form, we switch to the conformal momentum of the particle $\boldsymbol{p}=\boldsymbol{k}-e\boldsymbol{A}$ and introduce the conformal energy $\epsilon_{\boldsymbol{p}}\equiv\sqrt{\boldsymbol{p}^{2}+m^{2}a^{2}}$.  [^2] The quantities $\boldsymbol{p}$ and $\epsilon_{\boldsymbol{p}}$ do not coincide with the physical momentum and energy of the scalar particle measured by the comoving cosmological observer. They are introduced for convenience in further analysis. The corresponding physical quantities can be expressed as $\boldsymbol{p}_{\mathrm{phys}}=\boldsymbol{p}/a$ and $\epsilon_{\boldsymbol{p},\mathrm{phys}}=\epsilon_{\boldsymbol{p}}/a$. Further, we perform the following replacements in the system of equations (2):
@@
$\mathcal{F}_{\boldsymbol{k}}(\eta)\,\to\,\frac{d}{d\eta}\mathcal{F}(\eta,\boldsymbol{p})=\left[\frac{\partial}{\partial\eta}+\frac{\partial\boldsymbol{p}}{\partial\eta}\!\cdot\!\frac{\partial}{\partial\boldsymbol{p}}\right]\!\mathcal{F}(\eta,\boldsymbol{p})=\left[\frac{\partial}{\partial\eta}+e\,a^{2}\boldsymbol{E}\!\cdot\!\frac{\partial}{\partial\boldsymbol{p}}\right]\mathcal{F}(\eta,\boldsymbol{p})\,$ (23)
and similar relations also hold for the quantities $\mathcal{G}(\eta,\boldsymbol{p}),\mathcal{H}(\eta,\boldsymbol{p})$. Finally, we obtain the system of quantum Vlasov equations for the kinetic functions (21):
@@
These equations describe the creation of charged scalar particles in the expanding universe in the presence of spatially homogeneous electric field.

### v3

@@ Abstract
We study the Schwinger pair creation of scalar charged particles by a homogeneous electric field in an expanding universe in the quantum kinetic approach. We introduce an adiabatic vacuum for the scalar field based on the Wentzel–Kramers–Brillouin solution to the mode equation in conformal time and apply the formalism of Bogolyubov coefficients to derive a system of quantum Vlasov equations for three real kinetic functions. Compared to the analogous system of equations previously reported in the literature, the new one has two advantages. First, its solutions exhibit a faster decrease at large momenta which makes it more suitable for numerical computations. Second, it predicts no particle creation in the case of conformally coupled massless scalar field in the vanishing electric field, i.e., it respects the conformal symmetry of the system. We identify the ultraviolet divergences in the electric current and energy–momentum tensor of produced particles and introduce the corresponding counterterms in order to cancel them.
@@ 1 Introduction
One of the most fascinating phenomena occurring in a strong electromagnetic field is the Schwinger effect or the creation of particle–antiparticle pairs from physical vacuum [1, 2, 3]. It was predicted almost a hundred years ago and until the present day has not been observed in the laboratory because of extremely high value of the required electric field $E\sim 10^{18}\,\text{V}/\text{m}$ [4]. Nevertheless, such enormous fields might exist in the universe, e.g., around the compact objects like neutron stars or black holes [5, 6] or during the early stages of evolution (see, e.g., Refs. [7, 8, 9]).
The Schwinger effect in the early universe is often considered within the context of inflationary magnetogenesis models which include the coupling of the electromagnetic field to the inflaton or to the spacetime curvature [10, 11, 12]. The simplest scenario for the Schwinger pair production during inflation which can be studied analytically is the case of a constant and homogeneous electric field (with possible presence of a collinear magnetic field) in de Sitter spacetime which has been studied in much detail in the literature [7, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22]. The results of this simple approach has been used for a lot of phenomenological applications [23, 8, 24, 25, 26, 27, 28, 29, 9, 30, 31, 32, 33]; however, in realistic inflationary models both underlying assumptions, the constant in time gauge field and purely de Sitter expansion of the universe, typically do not hold.
In a time-dependent electromagnetic field, the Schwinger pair production has a more complicated dynamics. This was demonstrated by using the kinetic approach both in Minkowski space [34, 35, 36, 37, 38, 39, 40, 41, 42] and in the expanding universe [43, 44]. In contrast to the case of constant field, the Schwinger induced current exhibits the non-Markovian character depending not only on the field at a given moment of time but also on the prehistory. Moreover, due to inertial properties of charge carriers the current is retarded with respect to the changes in the electric field which leads to the oscillatory behavior of the induced current and the electric field.
In the kinetic approach, the particle creation is described by the Schwinger source term which can be constructed phenomenologically [34, 35, 43] or derived from the first principles [36, 37, 38, 39, 40, 41, 42, 44]. The latter way, although being much more involved, allows to capture the effects of particle statistics and the nonlocal in time character of the pair-creation process. There is, however, a fundamental issue which makes the application of such a first-principles approach ambiguous in the case of the expanding universe. This is a well-known in the literature problem of defining the vacuum state and particles in nonstationary backgrounds which do not match the Minkowski spacetime in the asymptotic past and future [45, 46, 47, 48]. The result for the number of produced particles depends on the definition of physical vacuum state or, in other words, on the choice of the observer.
In this work, we revisit the problem of pair creation of the scalar particles by a time-dependent electric field in the expanding Friedmann–Lemaître–Robertson–Walker (FLRW) universe previously considered in Ref. [44]. We mostly follow the same strategy in order to derive the system of quantum kinetic equations describing the Schwinger pair production in this system. However, in the present study, we perform all computations in terms of the conformal time rather then physical time considered in Ref. [44]. Despite the seemingly trivial change, we find that the definition of the adiabatic vacuum is different compared to the previous work and, consequently, the final system of equations appears to be different. We analyze the ultraviolet (UV) behavior of its solutions and show that kinetic functions exhibit faster decrease at large momenta as compared to the system in Ref. [44] which makes the new system of equations more attractive for numerical analysis. Also the new setup respects the conformal symmetry of the problem and does not lead to the particle production in the case of conformally coupled massless scalar field in the absence of the electric field.
The aim of this work is not to show that the computations in Ref. [44] are incorrect, but to demonstrate that there is a different and more convenient way to derive the system of quantum kinetic equations. Although we expect that the numerical result in the case of a strong electric field $|e\boldsymbol{E}|\gg H^{2},\,m^{2}$ (where $e$ is the electric charge, $H$ is the Hubble parameter, and $m$ is the particle’s mass) would be the same for both approaches, there might be some numerical differences in the case of a weak field where the particle production by the electric field and the time-varying metric are comparable.
The rest of the article has the following structure. In Sec. 2 we consider the evolution of a quantum scalar field on a classical background of the electric field and the expanding universe. In particular, we define the adiabatic vacuum and use the formalism of the Bogolyubov coefficients in order to derive the system of quantum kinetic equations. Then, in Sec. 3 we determine the basic observables—electric current and energy–momentum tensor of the produced particles—and express them in terms of the kinetic functions. In Sec. 4, we realize the renormalization program in order to separate the UV divergent contributions to the observables and cancel them by introducing the counterterms to the action. Section 5 is devoted to conclusions. In Appendix A, we determine the asymptotical behavior of the kinetic functions in the limit of large momenta.
Throughout the work, we use natural units and set $\hbar=c=1$. We assume that the universe is described by a spatially flat FLRW metric. In terms of conformal time $\eta$ it has the form $g_{\mu\nu}=\operatorname{diag}(a^{2},\,-a^{2},\,-a^{2},\,-a^{2})$, where $a=a(\eta)$ is the scale factor.
@@ 2 Quantum Vlasov equations for the scalar field
where the term $\mathcal{L}_{\mathrm{int}}$ describes the coupling of the electromagnetic field to some generic field $\phi$. Although this interaction is necessary to produce sufficiently strong electromagnetic field needed for the Schwinger pair production, in this work we will not stick to any specific model and will assume that the electromagnetic field already exists in the universe. Moreover, we consider the following configuration of the gauge field: (i) the magnetic field $\boldsymbol{B}$ is negligibly small compared to the electric one $\boldsymbol{E}$; (ii) the coherence length of the electric field $\lambda_{E}$ is much greater than any other physically relevant length scale in the problem, e.g., the Schwinger pair production scale $l_{\mathrm{S}}\sim|e\boldsymbol{E}|^{-1/2}$ or the Hubble scale $l_{H}\sim H^{-1}$. The latter condition allows us to treat the electric field spatially homogeneous and dependent only on time $\boldsymbol{E}=\boldsymbol{E}(\eta)$. Such a field configuration can be described by the vector potential (in the Coulomb gauge) $A_{\mu}=(0,\,\boldsymbol{A}(\eta))$ with $\boldsymbol{A}^{\prime}(\eta)=-a^{2}\boldsymbol{E}$. With the factor $a^{2}$, the electric field $\boldsymbol{E}$ is a physical field measured by the comoving observer.
Finally, we go back to the last term in Eq. (1) which represents the Lagrangian density of the complex scalar charged field $\chi$ with mass $m$ and charge $e$:
@@
Here $\lvert\chi\rvert^{2}\equiv\chi^{\dagger}\chi$, $\mathcal{D}_{\mu}=\partial_{\mu}-ieA_{\mu}$ is the covariant derivative acting on the scalar field, and $\xi$ is the coupling constant responsible for the nonminimal coupling of the scalar field $\chi$ to gravity.
Varying action in Eq. (1) with respect to $\chi^{\dagger}$, we obtain the equation of motion for the scalar field $\chi$:
@@
The factor $a(\eta)$ in denominator in Eq. (6) was introduced for further convenience.
Inserting the expansion (6) into the equation of motion (5) we get the mode equation governing the evolution of the mode function $\chi_{\boldsymbol{k}}(\eta)$:
@@
It has an oscillatorlike form with the time-dependent effective frequency
@@
It is easy to observe from Eq. (9) that in the absence of an electromagnetic field in a nonexpanding universe, the frequency $\Omega_{\boldsymbol{k}}$ does not depend on time $\eta$. Then, Eq. (8) has two independent solutions which describe the positive- and negative-frequency modes. If the initial condition imposed at a certain moment of time corresponds, e.g., to a positive-frequency mode, the solution remains to be positive-frequency all the time. Therefore, we conclude that the creation of particles in Minkowski spacetime does not occur.
Moreover, even in the expanding universe, for the massless scalar particle, $m=0$, with the conformal nonminimal coupling to gravity, $\xi=1/6$, and in the absence of the electric field, equation of motion (8) has a very simple oscillatorlike form $\chi^{\prime\prime}_{\boldsymbol{k}}+k^{2}\chi_{\boldsymbol{k}}=0$ whose positive-frequency solution is a well-known Bunch–Davies vacuum [49]:
@@
Being prepared in this state initially, the scalar field will remain in it forever meaning that there is no particle production even in the expanding universe. This is a consequence of the conformal invariance of the action of a massless conformally-coupled scalar field [48]. Here, we would like to note that this conformal symmetry becomes explicit only if we are working in the FLRW metric written in terms of the conformal time which is conformally flat.
The situation changes in the presence of a nonzero electric field in the expanding universe. Indeed, the value of $\Omega_{\boldsymbol{k}}$ becomes dependent on the conformal time $\eta$, and therefore, in general, it is impossible to find exact general solution to the mode equation (8) and separate positive- and negative-frequency modes [^2] Note, that in certain particular cases when the time dependences of the scale factor and the electric field have a simple form, one can find exact analytical solution to Eq. (8); see, e.g., Refs. [7, 13, 14] for the case of a minimally coupled complex scalar field in the constant electric field in de Sitter spacetime.. Nevertheless, one can still construct an approximate solution by employing the Wentzel–Kramers–Brillouin approximation up to a certain adiabatic order. For example, in the zeroth adiabatic order, the positive-frequency solution has the form
@@
and at the moment of time $\eta_{0}$ they satisfy the initial conditions $\alpha_{\boldsymbol{k}}(\eta_{0})=1$ and $\beta_{\boldsymbol{k}}(\eta_{0})=0$ (by construction of the adiabatic vacuum at the moment $\eta_{0}$).
It is straightforward to show that the mode equation (8) is identically satisfied if the evolution of the Bogolyubov coefficients is determined by the following system of equations:
@@
The main advantages of this system of equations compared to (15) is that it is real and does not contain fast oscillating coefficients. In order to rewrite it in the final form, we switch from the canonical momentum $\boldsymbol{k}$ to the kinetic momentum in the conformal spacetime $\boldsymbol{p}=\boldsymbol{k}-e\boldsymbol{A}$ and introduce the corresponding energy $\epsilon_{\boldsymbol{p}}\equiv\sqrt{\boldsymbol{p}^{2}+m^{2}a^{2}}$.  [^3] The quantities $\boldsymbol{p}$ and $\epsilon_{\boldsymbol{p}}$ do not coincide with the physical kinetic momentum and energy of the scalar particle measured by the comoving cosmological observer. They are introduced for convenience in further analysis. The corresponding physical quantities can be expressed as $\boldsymbol{p}_{\mathrm{phys}}=\boldsymbol{p}/a$ and $\epsilon_{\boldsymbol{p},\mathrm{phys}}=\epsilon_{\boldsymbol{p}}/a$. Further, we perform the following replacements in the system of equations (19):
@@
$\mathcal{F}^{\prime}_{\boldsymbol{k}}(\eta)\,\to\,\frac{d}{d\eta}\mathcal{F}(\eta,\boldsymbol{p})=\left[\frac{\partial}{\partial\eta}+\frac{\partial\boldsymbol{p}}{\partial\eta}\!\cdot\!\frac{\partial}{\partial\boldsymbol{p}}\right]\!\mathcal{F}(\eta,\boldsymbol{p})=\left[\frac{\partial}{\partial\eta}+e\,a^{2}\boldsymbol{E}\!\cdot\!\frac{\partial}{\partial\boldsymbol{p}}\right]\mathcal{F}(\eta,\boldsymbol{p})\,$ (23)
and similar relations also hold for the quantities $\mathcal{G}(\eta,\boldsymbol{p}),\mathcal{H}(\eta,\boldsymbol{p})$.
Finally, we obtain the system of quantum Vlasov equations for the kinetic functions (21):
@@
These equations describe the creation of charged scalar particles in the expanding universe in the presence of the spatially homogeneous electric field. We would like to mention an important property of this system of equations: it ensures that the combination $(1+2\mathcal{F})^{2}-(2\mathcal{G})^{2}-(2\mathcal{H})^{2}$ is conserved in time and equals to unity. This can also be seen directly from the definitions in Eqs. (16)–(18) together with normalization condition (14). According to Ref. [41], this is a manifestation of the charge conservation and is related to the invariance of the Casimir operator for the SU(1,1) algebra [50].